OpenStudy (anonymous):

Express the required measurement in exact form. What is the radius of a circle in which a 30° arc is 2 PI inches long? r = ? inches

OpenStudy (anonymous):

Ok, same problem, other way round number of arcs in a full circle: 360 / 30 = 12 circumference of that circle is hence: 12 * 2pi = 24pi 2*pi*r = 24pi r = 24pi / 2pi = 12 inches

OpenStudy (anonymous):

thanks alot